# Dependence of the Reconstructed Core-Collapse Supernova Gravitational Wave High-Frequency Feature on the Nuclear Equation of State, in Real Interferometric Data

### v1

@@ VI Conclusion
This paper shows the capability of using HFF slope measurements to distinguish between the possible EOS at play in CCSN events within 1 kpc for current-generation detectors. The improved sensitivity of future detectors will improve our ability to distinguish between different EOS, at larger distances. These improved sensitivities will also allow for an estimation of the second-order time dependence of the HFF, as the signal above 1 kHz will become detectable. We leave determinations of second-order time effects for future work. Other intrinsic properties of the CCSN progenitor that effect the HFF are also worth investigating in isolation, as was done here, and together. Connecting the HFF to intrinsic properties of the PNS and resulting neutron star is also a clear future goal. Measuring the time-dependent properties of the HFF would then directly relate to the evolution of the PNS mass and radius.
@@ Appendix A Curvature of the HFF
Appendix B Physical Interpretation of the HFF
In previous analyses of CCSN GW signals, the HFF is commonly identified as either a $g$-mode oscillation of the PNS – e.g. Kuroda et al. [27] – or a combination of $g$-mode and $f$-mode PNS oscillations – e.g. Torres-Forné et al. [71] and Morozova et al. [21]. These studies show correlation between the evolution in time of the HFF and the evolution in time of the frequency of the PNS quasi-normal modes. It is well known that the PNS $f$-mode will evolve in proportion to the mean density of the PNS,
$f_{f-\rm mode}\sim\sqrt{\frac{M_{\rm PNS}}{R_{\rm PNS}^{3}}},$ (8)
while the $g$-mode will evolve in proportion to the surface gravity of the PNS,
$f_{g-\rm mode}\sim\frac{GM_{\rm PNS}}{R_{\rm PNS}^{2}}.$ (9)
If the HFF is assumed to be some combination of $f$-mode and $g$-mode oscillations, the HFF time evolution can be parameterized by a combination of the mass and radius of the PNS – e.g., $M/R^{2}$ or $M/R^{3}$. These parameterizations are independent of EOS, but effects of the EOS may cause differences in the evolution of the PNS mass and radius.
It should be noted that studies use different definitions of $R_{\rm PNS}$ in Equations (8) and (9). In Müller et al. [18], an equation for the dominate GW frequency is derived assuming a surface $g$-mode oscillation, and they find good agreement with GW signals produced from two-dimensional CCSN simulations defining $R_{\rm PNS}$ as the $\mathrm{10^{11}}$$\mathrm{cm^{-3}}$ density contour. Torres-Forné et al. [22] use $R_{\rm shock}$ in place of $R_{\rm PNS}$ to parameterize their $f$-mode oscillations. Morozova et al. [21] define $R_{\rm PNS}$ as the radius at which the density is $\mathrm{10^{10}}$$\mathrm{cm^{-3}}$. Mezzacappa et al. [54] showed that the “universal” relation derived by Müller et al. [18] matched the GW peak frequency evolution best (for GW signals produced from three-dimensional CCSN simulations) at $R$ between the density contours of $\mathrm{10^{12}}$-$\mathrm{10^{13}}$$\mathrm{cm^{-3}}$, indicating that the assumption that the HFF is due to surface $g$-mode oscillations in these models is likely incorrect. Further, Rodriguez et al. [72] showed that mode-classification for the PNS is non-trivial, and multiple classification techniques may be necessary to gain a full understanding of the complex dynamics occurring in the evolving PNS.
In this preliminary exploration, we abstain from specifying the HFF as a particular mode and examine instead possible relations for the $f$-mode– and $g$-mode–dominated cases. As shown in Equations 8 and 9, the behavior of the HFF will be different based on whether it is due to an $f$-mode or a $g$-mode. Taking partial derivatives with respect to time of Equations 8 and 9, we find that the slope of the HFF is approximating either
$\frac{\partial f}{\partial t}\left|{}_{f\mathrm{-mode}}\sim\frac{1}{2}\right.f_{\mathrm{f-mode}}\left(\frac{\dot{M}}{M}-3\frac{\dot{R}}{R}\right)$ (10)
or
$\frac{\partial f}{\partial t}\left|{}_{g\mathrm{-mode}}\sim f_{\mathrm{g-mode}}\left(\frac{\dot{M}}{M}-2\frac{\dot{R}}{R}\right)\right.,$ (11)
at different times of the simulation. This is distinct from the parameterizations based on PNS properties of [71] and [25], though, ultimately, they should be related. We see that the slope of the HFF involves the evolution of the mass and radius of the oscillation region, with different factors multiplying the radius evolution term. While the definition of the oscillation region – e.g. whether the region is defined as bounded by the shock or by the PNS surface and, in the latter case, by how the PNS surface is defined – will affect the mass and radius evolution profiles, it is possible, for consistent definitions, that the HFF slope could be used to determine which mode is responsible for the HFF. Measuring the slope of the HFF thus has the potential to investigate which mode is dominant in the GW signal and how best to define the regional boundaries.

### v2

@@ VI Conclusion
This paper shows the capability of using HFF slope measurements to distinguish between the possible EOS at play in CCSN events within 1 kpc for current-generation detectors. The improved sensitivity of future detectors will improve our ability to distinguish between different EOS, for distances at least an order of magnitude larger. We also expect further benefits from customized tuning of cWB for PE purposes, as shown in Lin et al. [7]. These improved sensitivities will also allow for an estimation of the nonlinear time dependence of the HFF, given the signal above 1 kHz will become detectable. We leave a study of this phase of the evolution of the HFF to future work.
In this study, we investigated the effects on the HFF of varying the EOS in isolation, without varying progenitor parameters. Not only do parameters like rotation rate [48, 49, 50, 51, 52] and mass [53, 54, 55] affect the slope of the HFF, but parameters varied simultaneously affect the slope in more complex ways. As seen in Wang and Pan [56], varying the internal structure at the same time as rotation rate leads to HFF slopes that are decreased less than when there is no internal structure variation. Similarly, Jardine et al. [57] show that the presence of strong magnetic fields may also counteract the effect of rotation on the slope of the HFF.
Future studies investigating the combined effects of varying intrinsic properties of the progenitor and the EOS are necessary in order to develop a complete understanding of what can be learned about the EOS from the detection of the HFF in the next Galactic CCSN. The study conducted here can be viewed as providing limits on what we can hope to learn.